A sample of magnesium is heated and the mass of magnesium oxide produced found to be 80g. Calculate the mass of decomposed magnesium carbonate if the yield of the reaction is 80%.


1
Expert's answer
2021-08-18T05:32:01-0400

MgCO3 → CO2 + MgO

M(MgO) = 40.30 g/mol

n(MgO) "= \frac{80}{40.30}=1.985 \;mol"

Proportion:

1.985 – 80%

x – 100%

"x= \frac{1.985 \times 100}{80}= 2.481 \;mol"

According to the reaction:

n(MgCO3) = n(MgO)=2.481 mol

M(MgCO3) = 84.31 g/mol

m(MgCO3) "= 84.31 \times 2.481 = 209.20 \;g"

Answer: 209.20 g